OpenStudy (anonymous):

integrate sec^4x sqrt(1+tanx) dx

OpenStudy (anonymous):

hope it is something you are looking for

OpenStudy (anonymous):

wolfram doesn't help me understand the problem please explain the steps

OpenStudy (anonymous):

click show steps.. there are 2 times of substitute

OpenStudy (anonymous):

i want to view another way of doing it and not use 2 sub

OpenStudy (anonymous):

can you continue please

OpenStudy (anonymous):

i just stucked give me few mins

OpenStudy (anonymous):

okay no problem thank you

OpenStudy (anonymous):

oh sorry let tanx +1 = t secx^2dx = dt (1+(t-1)^2)t^1/2 (1+t^2+1-2t)t^1/2 {2t^1/2+t^5/2-2t^3/2}dt 4/3t^3/2+2/7t^7/2-4t^5/2 +c now put t = tanx + 1 you will get answer

OpenStudy (anonymous):

you need to draw this out, there are too many numbers i am sorry

OpenStudy (anonymous):

how can (1+(t-1)^2)t^1/2 be the next step if you subsituted wouldnt it be t^2 sqrt(t)

OpenStudy (anonymous):

check attachment

OpenStudy (anonymous):

oh i see what you did, but i was wondering how come the 1+tan^2x cannot become U^2 as well because U=1+tan^2x

OpenStudy (anonymous):

sorry but i didn't get it what u said

OpenStudy (anonymous):

i let 1+tanx = t

OpenStudy (anonymous):

nvm, i got it thank you

OpenStudy (anonymous):

welcome